OpenStudy (anonymous):

Manny loves to eat Cajun pizza. Each square inch of pizza is 8 calories. He orders a slice of pizza from a pizza with a diameter of 16 inches that is cut with a central angle of 20 degrees. How many calories does he eat?

OpenStudy (anonymous):

you'll need to find the area of the slice that has a central angle of 20 degrees... the total area of the the 16 inch pizza is A=pi*(16)^2 = 256pi sq inches. 20/360=1/18 of the whole pizza. so to find the amount of calories in 1/18 of the pizza, calculate 256*pi * 8 * (1/18)